Explain the term price elasticity of demand? How is it measured? What factors influence market

demand for products? If the price elasticity is -3 and RM 100 is the marginal cost of product

X, what should be the optimal sale price?

(Hint: apply the mark-up rule) 

1
Expert's answer
2021-12-30T12:25:37-0500

Price elasticity of demand refers to the degree of responsiveness of the quantity of goods or services demanded by the customers as a result of changes in price.

Price elasticity of demand can be measured by taking the percentage change in quantity demanded divided by the percentage change in price. Also, using mid-point method, and arc elasticity.

It is influenced by factors such as; price of related products such as substitutes or complements, the tastes and preferences of the consumer, income level of the consumers, and population of the market among other factors.

The mark up rule of pricing states that mark up "= (P-MC)\/P = 1\/Elasticity."

"-1\/3=(P-100)P"

"-P=3P-300"

"-4P=-300"

"P=75"

The Optimal Sale Price is RM 75